If you needed to collect 263 g of NaCl salt by evaporating off the water from a 4.2% by mass solution of NaCl salt water, how many grams of the solution do you need to evaporate the water off of?




1
Expert's answer
2021-02-18T07:46:39-0500

We have given,


Mass of "\text{NaCl}=236g"

Mass percent"=4.2" %


Let the mass of solution be "x"

We know,


Mass%="\dfrac{(\text{Mass of NaCl)}}{\text{(Mass of solution)}}\times 100"


"4.2=\dfrac{(263\times 100)}{x}"


"\implies x=\dfrac{26300}{4.2}\\\implies

x=6261.90g"


Hence, "6261.90g" of solution we need to evaporate the water off.